OpenStudy (camerondoherty):

Paul teaches the guitar in groups of three students. He rents a room for the guitar classes at $250 per month. He charges $30 per hour per student. Paul teaches six groups of students, and his net earnings per month are $1,910. Part A: If each student learns guitar for the same number of hours, write an equation that can be used to calculate the number of hours each student learns guitar in a month. (5 points) Part B: How many hours did each student learn guitar each month? Show your work and justify each step of your work. (5 points)

OpenStudy (anonymous):

Hm, this is a tricky one. First lets find his total earnings which we can find by adding 250m to the left side of the equation, where m is the months since he is charged 250 per month. So far we have: \(250m+1910=\) Now for the students part. It says he teaches 3 students at a time, where it costs 30dollars an hour per student. So that is the same as \((3*30h)\), but we are not done yet because he teaches 6 of these groups so we have to multiply that by 6. This gives us \(6(3*30h)\) so the equation is \(250m+1910=6(90h)\) which is the same as \(250m+1910=540h\)

OpenStudy (anonymous):

To solve the second part you plug in 1 for m which means for 1 month and then solve for what h=

OpenStudy (camerondoherty):

Thank You @Venny